OpenStudy (anonymous):

f(x)= 3x+1, then f(a+h)- f(a)

hero (hero):

Hint: f(x) = 3x + 1 f(a) = 3a + 1 f(a + h) = 3(a + h) + 1

OpenStudy (anonymous):

so you would distribute 3(a+h)

hero (hero):

First see if you can figure out what f(a + h) - f(a) is

OpenStudy (anonymous):

hmm im not sure how to solve that either

hero (hero):

f(a + h) - f(a) = 3(a +h) + 1 - (3a + 1)

hero (hero):

Simplify that